OpenStudy (anonymous):

step by step simplify this expression (3+i)-(7+6i)

OpenStudy (anonymous):

same as combine like terms, as if you had \((3+x)-(7+6x)\)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Well you can open the parenthesis, which gives \[(3+i)-(7+6i) = 3+i-7-6i = -4-5i\]

OpenStudy (anonymous):